Multiple-Product Breakeven

Parker Pottery produces a line of vases and a line of ceramic figurines. Each line uses the same equipment and labor; hence, there are no traceable fixed costs. Common fixed cost equals $33,600. Parker's accountant has begun to assess the profitability of the two lines and has gathered the following data for last year:

Vases Figurines
Price $40 $70
Variable cost 30 42
Contribution margin $10 $28
Number of units 1,000 500

Required:

If required, round your final answers to nearest whole value.

1. Compute the number of vases and the number of figurines that must be sold for the company to break even.

Break-even vases units
Break-even figurines units

2. Parker Pottery is considering upgrading its factory to improve the quality of its products. The upgrade will add $5,280 per year to total fixed cost. If the upgrade is successful, the projected sales of vases will be 2,000, and figurine sales will increase to 1,000 units. What is the new break-even point in units for each of the products?

Break-even vases units
Break-even figurines units
Answer #1

1) Weighted average contribution margin per unit = (10*2/3+28*1/3) = 16 per unit

Total break even point = 33600/16 = 2100 Units

Break even vases = 2100*2/3 = 1400 Units

Break even figurines = 2100*1/3 = 700 Units

2) Total break even point = (33600+5280)/16 = 2430 Units

Break even vases = 2430*2/3 = 1620 Units

Break even figurines = 2430*1/3 = 810 Units